# Image Corruption-Inspired Membership Inference Attacks against Large Vision-Language Models

### v2

@@ Abstract
Large vision-language models (LVLMs) have demonstrated outstanding performance in many downstream tasks. However, LVLMs are trained on large-scale datasets, which can pose privacy risks if training images contain sensitive information. Therefore, it is important to detect whether an image is used to train the LVLM. Recent studies have investigated membership inference attacks (MIAs) against LVLMs, including detecting image-text pairs and single-modality content. In this work, we focus on detecting whether a target image is used to train the target LVLM. We design simple yet effective Image Corruption-Inspired Membership Inference Attacks (ICIMIA) against LLVLMs, which are inspired by LVLM’s different sensitivity to image corruption for member and non-member images. We first perform an MIA method under the white-box setting, where we can obtain the embeddings of the image through the vision part of the target LVLM. The attacks are based on the embedding similarity between the image and its corrupted version. We further explore a more practical scenario where we have no knowledge about target LVLMs and we can only query the target LVLMs with an image and a question. We then conduct the attack by utilizing the output text embeddings’ similarity. Experiments on existing datasets validate the effectiveness of our proposed attack methods under those two different settings.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) (Liu et al., 2024a, b; Bai et al., 2023b; Achiam et al., 2023), which can generate text outputs based on visual and/or textual input, are attracting increasing attention. Many LVLMs have been developed, which have shown great performance on various tasks (Dong et al., 2024; Xu et al., 2024; Wu et al., 2025; Yue et al., 2024; Li et al., 2024a; Bucciarelli et al., 2024) such as biomedical question answering (Li et al., 2023a).
Figure 1: An example of some potential risks in LVLMs.
@@
Despite their great success, LVLMs have also raised critical concerns about privacy and copyright issues as shown in Figure 1. LVLMs are usually trained on large-scale datasets (Ng et al., 2020; Changpinyo et al., 2021; Byeon et al., 2022). The training data might contain sensitive information, such as unauthorized medical data and copyrighted content. Previous works have shown that neural networks, especially large models, can memorize training data (Song et al., 2017; Carlini et al., 2019, 2022b). Thus, the memorization phenomena might also happen in LVLMs and inadvertently cause training data leakage (Li et al., 2024b), which could cause substantial loss to data owners. Hence, knowing whether one’s data is used to train an LVLM is important for privacy and copyright protection.
@@
Membership inference attacks (MIAs), which aim to determine whether a given sample is used to train a model (Hu et al., 2022a; Shokri et al., 2017), are critical for ensuring data integrity (Oren et al., 2023; Duan et al., 2024). Generally, models tend to overfit to the training data, resulting in higher prediction confidence for member data (data used for training) than non-member data. Many traditional MIA methods adopt such nuance difference in model prediction to differentiate members and non-members.
Recently, some works (Ko et al., 2023; Li et al., 2024b; Hu et al., 2025) have explored the MIA on vision-language models, from CLIP (Radford et al., 2021) to LVLMs (Zhang et al., 2023). Ko et al. (2023) focus on determining whether an image-text pair exists in the training data of CLIP. However, detecting whether an image in the training data is more practical than detecting the entire image-text pair as the image owner might only have the image without text while the text description might be labeled by the model trainer (Li et al., 2024b).
@@
Therefore, we study the problem of single-image based MIA against LVLMs. The work in this direction is rather limited Li et al. (2024b). Li et al. (2024b) also conduct MIA on a single modality (Image or textual description) against LVLMs by using the output logits of LVLM. Instead of using output logits, we propose a new perspective, i.e., adopting image embedding from the visual part of LVLMs. Our intuition is: as the model has seen member images, it should be able to give robust image embedding of a member image even if some details of the image are missing. In other words, the image embedding of a member is more robust to image corruption than that of non-member images, which is verified by our preliminary experiment in Section 3 (see Fig. 3). Based on this observation, we propose a novel MIA algorithm under the white-box setting, where we can obtain the image embedding from LVLMs. Given an image, we corrupt the image and utilize the image embedding similarity between the raw image and the corrupted version to determine if an image is a member. A higher similarity means the image is more likely to be a member.
@@
Our main contributions are: (i) In this work, we investigate two membership inference attack scenarios targeting LVLMs. For each setting, we propose one simple yet strong attack method that leverages the model’s robustness to image corruption on its training images; (ii) Extensive experiments on existing datasets show the effectiveness of the proposed method.
2 Related Works
@@ 2.1 Large Vision-Language Models
Large Vision-Language Models (Bai et al., 2025; Chen et al., 2024a, b; Liu et al., 2024b, a; Chen et al., 2024c; Tong et al., 2024; Zhu et al., 2025; Young et al., 2024; Zhu et al., 2023; Chen et al., 2023; Li et al., 2024c; Du et al., 2025), also known as multimodal large language models (Fu et al., 2024) are developing rapidly due to the success of language models (Wang et al., 2024; Zhao et al., 2023; Chiang et al., 2023; Touvron et al., 2023a, b; Grattafiori et al., 2024; Yang et al., 2025, 2024; Jiang et al., 2023; Team, 2023; Penedo et al., 2023; Lin et al., 2024) such as Qwen (Bai et al., 2023a; Yang et al., 2024) and LLaMA series (Touvron et al., 2023a, b). These models, like LLaVA 1.5 (Liu et al., 2024c), can generate textual outputs given textual questions and images.
@@ 2.2 Membership Inference Attack
One main category of MIA methods is metric-based (Hu et al., 2022a; Li et al., 2024b; Sablayrolles et al., 2019; Choquette-Choo et al., 2021; Li et al., 2021), which use some well-designed metrics, such as the entropy of output logits, to determine the membership status of a given sample. Another main category is methods based on shadow training (Shokri et al., 2017), which trains some shadow models to simulate the target model and then conducts MIAs based on these shadow models. Many later works have started to explore the MIA on LLMs (Mattern et al., 2023; Mireshghallah et al., 2022; Ren et al., 2024; Shi et al., 2024; Zhang et al., 2024) such as Min-K(Shi et al., 2024) and Min-K%++ (Zhang et al., 2024).
With the development of multimodal learning, there are also some works exploring MIA on multimodal models (Hu et al., 2022b; Ko et al., 2023; Li et al., 2024b). We focus on vision-text here. Liu et al. (2021) studies MIAs on image encoder models such as CLIP vision encoder (Radford et al., 2021). It calculates the similarity scores between the augmented images’ embeddings and the scores are then used to train a classifier to infer the member status of an image. Our work is inspired by them. Ko et al. (2023) work on detecting whether an image-text pair is in the training data of CLIP models. Li et al. (2024b) investigates the single-modality MIA in LVLMs, which is a more practical scenario. They calculate the Rényi Entropy (Rényi, 1961) based on different slices of logits. Hu et al. (2025) found that member data and non-member data have different sensitivity to temperature. They then perform four attack methods under four different settings to detect whether a group of images is used in the training stage.
@@ 3 Preliminaries
In this section, we give the necessary background information and formulate the problems.
@@ 3.1 Large Vision-Language Models
As shown in Fig. 2, an LVLM $\mathbf{M}_{\theta}$ usually consists of three parts: A vision encoder $f_{Vision}$, an LLM $f_{LLM}$, and a modality connection module $f_{Align}$. The vision encoder $f_{Vision}$, e.g., CLIP Radford et al. (2021), takes an image $\mathbf{x}$ as input and outputs the embeddings of $N$ image patches $\{\mathbf{z}_{1},...,\mathbf{z}_{N}\}$, where $\mathbf{z}_{i}\in\mathbf{R}^{d_{v}}$ is the $i$-th patch embedding with $d_{v}$ being the embedding dimension. The patch embeddings are then transformed into an embedding sequence $\mathbf{E}^{v}=\{\mathbf{e}_{1}^{v},...,\mathbf{e}_{N}^{v}\}$ which is in the embedding space of $f_{LLM}$ using $f_{Align}$. Each element in $\mathbf{E}^{v}$ can be represented as:
@@
The text prompt $T_{in}$, composed of the instruction and the question, is encoded into an embedding sequence of tokens $\mathbf{E}^{t}=\{\mathbf{e}_{1}^{t},...,\mathbf{e}_{K}^{t}\}$, where $K$ is the number of tokens. Finally, the text embedding and the image embeddings are used as input to the LLM $f_{LLM}$ to get the text output as $T_{out}=f_{LLM}(\mathbf{E}^{v},\mathbf{E}^{t})$.
Figure 3: A histogram of similarity scores of corrupted images’ (Through Gaussian Blur) embeddings and original images’ embeddings for member and non-member data. Scores in this figure achieve an AUC of 0.881.
Figure 4: A histogram of similarity scores of textual output embeddings between corrupted images (Through Gaussian Blur) and the original images for member and non-member data. Scores in this figure achieve an AUC of 0.652.
@@ 3.2 Threat Model
Attacker’s Goal. Given a trained LVLM $\mathbf{M}_{\theta}$, the attacker’s goal is to determine if a specific image was in the training set of the target LVLM $\mathbf{M}_{\theta}$. The attackers only have the target image and do not require its corresponding ground-truth text description. The problem is formally defined as:
@@
Attacker’s Knowledge. In our work, we consider two practical settings, i.e., white-box and black-box. (i) White-box Setting. In this setting, the attacker can get the embedding of the image, i.e., the embedding through the model’s vision encoder and modality alignment module. This is practical for open-source LVLMs. (ii) Black-box Setting. In this setting, the attacker can only query the LVLM $\mathbf{M}_{\theta}$ with image input $\mathbf{x}$ and text input $T_{in}$ to get the text output $T_{out}$. The attacker has no knowledge of the target model. This setting is more realistic for commercial LVLMs.
@@ 3.3 Intuition for Performing MIA
As we are only given the image, how to fully utilize the image to determine if the image remains a challenging question. Inspired by the findings of Liu et al. (2021) that a CLIP vision encoder tends to overfit to its training data and member images show higher similarity for two augmented versions, we assume that the image embedding of a member image from the vision encoder and modality alignment module should be more robust to image corruption than that of non-member images. The intuition is, if the model has seen the image and memorized the image, then even if the image is corrupted, e.g., some details missing, the model can still recall the details and result in embedding similar to the original one. To verify our assumption, for each image, we first apply Gaussian Blur (Bradski, 2000) as the corruption method to obtain the corrupted image, where Gaussian Blur is a smoothing technique that reduces image detail and noise by averaging pixels with a weighted Gaussian kernel. We compute the similarity score between the embedding of the original image and that of its corrupted version. The chosen model is LLaVA-1.5-7B (Liu et al., 2024a), and the dataset is the Img_Flickr dataset constructed by Li et al. (2024b). The similarity scores of member images and non-member images are shown in Figure 3. We can observe that member images normally show a higher similarity score than non-member images, which aligns well with our assumption. The scores show a clear difference between the two groups, which suggests that the similarity score between the original image embedding and its corrupted version’s embedding can be used to perform an image membership inference attack. Thus, we can utilize the robustness of member image embedding to determine its membership.
However, in black-box setting, we are unable to obtain the image embedding. As image embedding of member image is robust, it might also result in robust text even if the image is corrupted. Thus, we assume that given the same text prompt (Instruction), an image seen during the LVLM’s training process will produce a text output that is more similar to the output generated from its corrupted version, compared to images not included in the training data. To verify our assumption, we use the same dataset, model, and corruption method as above. We compute the similarity score between the output text embeddings of the original and corrupted versions of each image, given the same prompt, separately. The results are shown in Figure 4. We observe that the discrepancy in similarity between member and non-member images still exists and can serve as a metric, although it is less apparent in Figure 3.
@@ 4.1 White-Box MIA via Image Similarity
6:Compute similarity score $s_{img}$ via Eq. 2
@@
We first discuss the grey-box setting where we can obtain the embeddings of a given image through the vision encoder and modality alignment module.
@@
where $\mathbf{e}_{i}^{v}$ is the embedding of the $i$-patch of image $\mathbf{x}_{t}$ obtained by Eq. 1 and $\mathbf{e}_{i}^{v^{\prime}}$ denotes the $i$-patch embedding of the corrupted image $\mathbf{x}_{t}^{\prime}$. $N$ is the number of patches. Sim is a similarity function. We use cosine similarity here. The larger the $s_{img}$ is, the more likely the image is viewed as a member image. The target image $\mathbf{x}_{t}$ is predicted as a member image if $s_{img}$ is bigger than a threshold $\lambda$. Algorithm 1 summarizes the image similarity-based attack.
@@ 4.2 Black-box MIA via Text Similarity
where $Emb()$ denotes a text embedding model, such as OpenAI’s text-embedding-3-small model [^1] https://platform.openai.com/docs/models/text-embedding-3-small. Sim is a similarity function as in Eq. 2 which is a cosine similarity function. Similarly, $x_{t}$ is considered member data if $s_{text}>\lambda$. This method is summarized in Algorithm 2. This attack is similar to the Target-only Inference in Hu et al. (2025) where they also use text similarity scores. The difference is that they use the average similarity score between text outputs generated by querying the model multiple times.
@@
7:Calculate the output similarity using Eq. 3
@@
5 Experiment
In this section, we evaluate our methods on existing datasets to answer the following research questions: (i) How well do our proposed methods perform in conducting membership inference attacks against large vision-language models? and (ii) What are the impacts of hyperparameters?

### v3

@@ Abstract
Large vision-language models (LVLMs) have demonstrated outstanding performance in many downstream tasks. However, LVLMs are trained on large-scale datasets, which can pose privacy risks if training images contain sensitive information. Therefore, it is important to detect whether an image is used to train the LVLM. Recent studies have investigated membership inference attacks (MIAs) against LVLMs, including detecting image-text pairs and single-modality content. In this work, we focus on detecting whether a target image is used to train the target LVLM. We design simple yet effective Image Corruption-Inspired Membership Inference Attacks (ICIMIA) against LVLMs, which are inspired by LVLM’s different sensitivity to image corruption for member and non-member images. We first perform an MIA method under the white-box setting, where we can obtain the embeddings of the image through the vision part of the target LVLM. The attacks are based on the embedding similarity between the image and its corrupted version. We further explore a more practical scenario where we have no knowledge about target LVLMs and we can only query the target LVLMs with an image and a textual instruction. We then conduct the attack by utilizing the output text embeddings’ similarity. Experiments on existing datasets validate the effectiveness of our proposed methods under those two different settings.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) (Liu et al., 2024a, b; Bai et al., 2023), which can generate text outputs based on visual and/or textual input, are attracting increasing attention. Many LVLMs have been developed, which have shown great performance on various tasks (Dong et al., 2024; Xu et al., 2024; Wu et al., 2025; Yue et al., 2024; Li et al., 2024a; Bucciarelli et al., 2024) such as biomedical question answering (Li et al., 2023a). Despite their great success, LVLMs have also raised critical concerns about privacy and copyright issues. LVLMs are usually trained on large-scale datasets (Changpinyo et al., 2021; Byeon et al., 2022). The training data might contain sensitive information, such as unauthorized medical data and copyrighted content. Previous works have shown that neural networks, especially those trained on large-scale datasets, can memorize training data (Song et al., 2017; Carlini et al., 2019, 2023). Thus, the memorization phenomena might also happen in LVLMs and inadvertently cause training data leakage (Li et al., 2024b), which could cause substantial loss to data owners. Hence, knowing whether one’s data is used to train an LVLM is important for privacy and copyright protection.
@@
Membership inference attacks (MIAs), which aim to determine whether a given sample is used to train a model (Hu et al., 2022a; Shokri et al., 2017), are critical for ensuring data safety and investigating data contamination (Oren et al., 2024; Duan et al., 2024). Generally, models tend to overfit to the training data, resulting in higher prediction confidence for member data (data used for training) than non-member data. Many traditional MIA methods (Shokri et al., 2017) adopt such nuance difference in model prediction to differentiate members and non-members. Recently, some works (Ko et al., 2023; Li et al., 2024b; Hu et al., 2025) have explored the MIA on vision-language models, from CLIP (Radford et al., 2021) to LVLMs (Zhang et al., 2023). Ko et al. (2023) focus on determining whether an image-text pair exists in the training data of CLIP. However, detecting whether an image in the training data is more practical than detecting the entire image-text pair (Li et al., 2024b) as the image owner might only have the image without text while the text description might be labeled by the model trainer.
@@
Therefore, we study the problem of single-image MIA against LVLMs. The work in this direction is rather limited Li et al. (2024b). Li et al. (2024b) also conduct MIA on a single modality (Image or textual description) against LVLMs by using the output logits of LVLM. Instead of using output logits, we propose a new perspective, i.e., adopting image embedding from the visual part of LVLMs. Our intuition is: as the model has seen member images, it should be able to give robust image embedding of a member image even if some details of the image are missing. In other words, the image embedding of a member is more robust to image corruption than that of non-member images, which is verified by our preliminary experiment in Section 3 (see Figure. 2). Based on this observation, we propose a novel MIA algorithm under the white-box setting, where we can obtain the image embedding from LVLMs. Given an image, we corrupt it and use the image embedding similarity between the raw image and the corrupted version to decide if it is a member. A higher similarity means the image is more likely to be a member.
@@
Our main contributions are: (i) In this work, we investigate two membership inference attack scenarios targeting LVLMs. For each setting, we propose one simple yet strong attack method that leverages the model’s robustness to image corruption on its training images; (ii) Extensive experiments on existing datasets show the effectiveness of the proposed membership inference method.
2 Related Work
@@ 2.1 Large Vision-Language Models
Large Vision-Language Models (Liu et al., 2024b, a; Chen et al., 2024b; Tong et al., 2024; Zhu et al., 2024; Chen et al., 2023; Li et al., 2024c; Chen et al., 2024a), also known as multimodal large language models (Fu et al., 2024) are developing rapidly due to the success of language models across various aspects (Wang et al., 2025; Zhao et al., 2023; Zhang et al., 2025b; Chiang et al., 2023; Touvron et al., 2023a, b; Team, 2023; Penedo et al., 2023; Lin et al., 2024). These multimodal models, like LLaVA series (Liu et al., 2023, 2024a), can generate textual outputs given textual questions and images.
@@ 2.2 Membership Inference Attack
One main category of MIA methods is metric-based (Sablayrolles et al., 2019; Choquette-Choo et al., 2021), which use some well-designed metrics (Hu et al., 2022a), such as the metrics based on loss, to determine the membership status of a given sample. Another main category consists of methods based on shadow training (Shokri et al., 2017), which trains some shadow models to simulate the target model and then conducts MIAs based on these shadow models. Many later works (Mireshghallah et al., 2022; Mattern et al., 2023; Ren et al., 2025; Shi et al., 2024), such as Min-K% (Shi et al., 2024) and Min-K%++ (Zhang et al., 2025a), have started to explore the MIA on Large Language Models (LLMs).
With the development of multimodal learning, there are also some works exploring MIA on multimodal models (Hu et al., 2022b; Ko et al., 2023; Hu et al., 2025; Li et al., 2024b). We focus on vision-text here. EncoderMI (Liu et al., 2021) studies MIAs on image encoder models such as CLIP vision encoder (Radford et al., 2021). It calculates the similarity scores between the augmented images’ embeddings and the scores are then used to train a classifier to infer the member status of an image. Our work is inspired by EncoderMI. Ko et al. (2023) work on detecting whether an image-text pair is in the training data of CLIP models. Li et al. (2024b) investigate the single-modality MIA in LVLMs, which is a more practical scenario. They calculate the Rényi Entropy (Rényi, 1961) based on different slices of logits to infer membership. Hu et al. (2025) found that member data and non-member data have different sensitivity to temperature. They then perform four attack methods under four different settings to detect whether images are used in LVLM’s training stage.
@@ 3 Preliminaries
In this section, we give the necessary background information and formulate the problems for MIAs against LVLMs.
@@ 3.1 Large Vision-Language Models
As shown in Figure 1, an LVLM $\mathbf{M}_{\theta}$ usually consists of three parts (Liu et al., 2024a): A vision encoder $f_{Vision}$, an LLM $f_{LLM}$, and a modality connection module $f_{Align}$. The vision encoder $f_{Vision}$, e.g., CLIP Radford et al. (2021) vision encoder used in LLaVA-1.5, takes an image $\mathbf{x}$ as input and outputs the embeddings of $N$ image patches (excluding the CLS token) $\{\mathbf{z}_{1},...,\mathbf{z}_{N}\}$, where $\mathbf{z}_{i}\in\mathbf{R}^{d_{v}}$ is the $i$-th patch embedding with $d_{v}$ being the embedding dimension. The patch embeddings are then transformed into an embedding sequence $\mathbf{E}^{v}=\{\mathbf{e}_{1}^{v},...,\mathbf{e}_{N}^{v}\}$ which is in the embedding space of $f_{LLM}$ using $f_{Align}$. Each element in $\mathbf{E}^{v}$ can be represented as:
@@
The text prompt $T_{in}$, composed of the instruction and the question, is tokenized and then encoded into an embedding sequence of tokens $\mathbf{E}^{t}=\{\mathbf{e}_{1}^{t},...,\mathbf{e}_{K}^{t}\}$, where $K$ is the number of tokens. Finally, the text embedding and the image embeddings are used as input to the LLM $f_{LLM}$ to get the text output $T_{out}$.
Figure 2: A histogram of similarity scores of corrupted images’ embeddings and original images’ embeddings for member and non-member data.
Figure 3: A histogram of similarity scores of textual output embeddings between corrupted images and the original images for member and non-member data.
@@ 3.2 Threat Model
Attacker’s Goal. Given a trained LVLM $\mathbf{M}_{\theta}$, the attacker’s goal is to determine if a specific image was in the training set of $\mathbf{M}_{\theta}$. The attackers only have the target image and do not require its corresponding ground-truth text description. The problem is formally defined as:
@@
Attacker’s Knowledge. In our work, we consider two practical settings, i.e., white-box and black-box: (i) White-box Setting. In this setting, the attacker can get the embedding of the image, i.e., the embedding through the model’s vision encoder and modality alignment module. This is practical for open-source LVLMs. $T_{in}$ is not required in this setting. (ii) Black-box Setting. In this setting, the attacker can only query the LVLM $\mathbf{M}_{\theta}$ with image input $\mathbf{x}$ and text input $T_{in}$ to get the text output $T_{out}$. The attacker has no knowledge of the target model. This setting is more realistic for commercial LVLMs.
@@ 3.3 Intuition for Performing MIA
As we are only given the image, how to fully utilize the image to determine if the image remains a challenging question. Inspired by the findings of EncoderMI (Liu et al., 2021) that a CLIP vision encoder tends to overfit to its training data and member images show higher similarity for two augmented versions, we explore whether LVLMs have similar features under image corruption methods. We assume that the image embedding of a member image from the vision encoder and modality alignment module should be more robust to image corruption than that of non-member images. The intuition is, if the model has seen the image and memorized the image, then even if the image is corrupted, e.g., some details missing, the model can still recall the details and result in embedding similar to the original one. To verify our assumption, for each image, we first apply Gaussian Blur as the corruption method to obtain the corrupted image, where Gaussian Blur is a smoothing technique that reduces image detail and noise by averaging pixels with a weighted Gaussian kernel. We compute the similarity score between the embedding of the original image and that of its corrupted version. The chosen model is LLaVA-1.5-7B (Liu et al., 2024a), and the dataset is the VL-MIA/Flickr dataset constructed by Li et al. (2024b). The similarity scores of member images and non-member images are shown in Figure 2. We can observe that member images normally show a higher similarity score than non-member images, which aligns well with our assumption. The scores show a clear difference between the two groups, which suggests that the similarity score between the original image embedding and its corrupted version’s embedding can be used to perform an image membership inference attack. Thus, we can utilize the robustness of member image embedding to determine its membership.
However, in the black-box setting, we are unable to obtain the image embedding. As image embedding of member image is robust, it might also result in robust text even if the image is corrupted. Thus, we assume that given the same text prompt (Instruction), an image seen during the LVLM’s training process will produce a text output that is more similar to the output generated from its corrupted version, compared to images not included in the training data. To verify our assumption, we use the same dataset, model, and corruption method as above. We compute the similarity score between the output text embeddings of the original and corrupted versions of each image, given the same prompt, separately. The results are shown in Figure 3. We observe that the discrepancy in similarity between member and non-member images still exists and can serve as a metric, although it is less apparent than that in Figure 2.
@@ 4.1 White-Box MIA via Image Similarity
6:Compute similarity score $s_{img}$ via Equation 2
@@
We first discuss the white-box setting where we can obtain the embeddings of a given image through the vision encoder and modality alignment module.
@@
where $\mathbf{e}_{i}^{v}$ is the embedding of the $i$-patch of image $\mathbf{x}_{t}$ obtained by Equation 1 and $\mathbf{e}_{i}^{v^{\prime}}$ denotes the $i$-patch embedding of the corrupted image $\mathbf{x}_{t}^{\prime}$. $N$ is the number of patches. $\text{Sim}(,)$ is a similarity function. We use cosine similarity here. The larger the $s_{img}$ is, the more likely the image is viewed as a member image. The target image $\mathbf{x}_{t}$ is predicted as a member image if $s_{img}$ is bigger than a threshold $\lambda$. Algorithm 1 summarizes the image similarity-based attack.
@@ 4.2 Black-box MIA via Text Similarity
where $Emb()$ denotes a text embedding model, such as OpenAI’s text-embedding-3-small model [^1] https://platform.openai.com/docs/models. $\text{Sim}(,)$ is a similarity function as in Equation 2 which is also a cosine similarity function. Similarly, $x_{t}$ is considered member data if $s_{text}>\lambda$. This method is summarized in Algorithm 2. This attack is similar to the Target-only Inference in Hu et al. (2025) where they also use text similarity scores. The difference is that they use the average similarity score between text outputs generated by querying the model multiple times.
@@
7:Calculate the output similarity using Equation 3
@@
5 Experiments
In this section, we evaluate our proposed ICIMIA on representative LVLM MIA datasets to answer these research questions: (i) How well do our proposed ICIMIA perform in conducting membership inference attacks against large vision-language models? and (ii) What are the impacts of hyperparameters?